OpenStudy (anonymous):

Choose the equation of the line passing through the point (4, -2) and perpendicular to y = x - 7. y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14

OpenStudy (shubhamsrg):

none i guess..

OpenStudy (campbell_st):

lol... freddie... maybe you miss typed the question

OpenStudy (anonymous):

is not

OpenStudy (anonymous):

Choose the equation of the line passing through the point (4, -2) and perpendicular to y =1/3x - 7. y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14

OpenStudy (anonymous):

y = -3x + 10